### v2

# ReCAP: Recursive Cross Attention Network for Pseudo-Label Generation in Robotic Surgical Skill Assessment

Abstract
In surgical skill assessment, Objective Structured Assessments of Technical Skills (OSATS scores) and the Global Rating Scale (GRS) are established tools for evaluating the performance of surgeons during training. These metrics, coupled with feedback on their performance, enable surgeons to improve and achieve standards of practice. Recent studies on the open-source dataset JIGSAW, which contains both GRS and OSATS labels, have focused on regressing GRS scores from kinematic signals, video data, or a combination of both. In this paper, we argue that regressing the GRS score, a unitless value, by itself is too restrictive, and variations throughout the surgical trial do not hold significant clinical meaning. To address this gap, we developed a recurrent transformer model that outputs the surgeon’s performance throughout their training session by relating the model’s hidden states to five OSATS scores derived from kinematic signals. These scores are averaged and aggregated to produce a GRS prediction, enabling assessment of the model’s performance against the state-of-the-art (SOTA). We report Spearman’s Correlation Coefficient (SCC), demonstrating that our model outperforms SOTA models for all tasks, except for Suturing under the leave-one-subject-out (LOSO) scheme (SCC 0.68-0.89), while achieving comparable performance for suturing and across tasks under the leave-one-user-out (LOUO) scheme (SCC 0.45-0.68) and beating SOTA for Needle Passing (0.69). We argue that relating final OSATS scores to short instances throughout a surgeon’s procedure is more clinically meaningful than a single GRS score. This approach also allows us to translate quantitative predictions into qualitative feedback, which is crucial for any automated surgical skill assessment pipeline. A senior surgeon validated our model’s behaviour and agreed with the semi-supervised predictions 77 % ($p=0.006$) of the time.
I Introduction
Skill assessment is the cornerstone of every surgical training regime. It allows trainers and trainees to quantify progress and performance across in surgery. Traditionally, this evaluation has relied on subjective feedback from experienced surgeons, resulting in variability across institutions and regions [4].
Recognizing the need for a standardised framework for surgical skill assessment, the Objective Structured Assessment of Technical Skills (OSATS) [19] score was developed. This Likert-style scale assesses various core components in surgery. The summation of individual OSATS scores yields a Global Rating Score (GRS), utilised to assess surgical trainees’ performance across multiple procedures. OSATS provides quantitative and qualitative performance analysis, irrespective of the level of expertise.
While these metrics are able to capture an overall picture of performance, they rely on the assessors’ feedback, typically senior clinicians [2], to capture intraoperative performance [3] during surgical trials.
Despite the presence of a standardised framework for surgical skill assessment, the process remains time-consuming and susceptible to rater bias. Moreover, the time constraints faced by senior clinicians significantly limit the training opportunities available to trainees outside the operating theatres [5]. The advent of machine learning (ML) and deep learning (DL) presents a promising avenue for automating the assessment process, thereby mitigating these challenges and expanding training access.
The JHU-ISI Gesture and Skill Assessment Working Set (JIGSAW) dataset [30] is currently used as the benchmark dataset for surgical skill assessment. While recent studies on the JIGSAW dataset have shifted towards video-centric approaches, the field of surgical skill assessment still relies on kinematic data [23, 25, 26, 24]. Kinematic data offer standardisation across datasets whereas video data is prone to considerable variations. Moreover, kinematic data imposes lower computational demands, allowing for rapid and efficient development processes. Lastly, we argue that the use of kinematic data instead of video is more ethical. Considering these factors, we direct our focus towards kinematic data.
While existing ML models excel at categorising surgical expertise [14, 8] and predicting overall performance metrics like the GRS  [17, 18, 20, 8, 14] on the JIGSAW dataset  [30], they often fall short in capturing nuanced variations throughout a surgical procedure. Recent research efforts have aimed to address this gap by exploring models that can track changes in GRS or OSATS scores over the course of a procedure [14, 22, 12, 20, 8, 13, 17]. Yet, many of these approaches either increase clinician workload [12, 20] by requiring additional labels or lack meaningful validation [8, 13, 17]. It’s also important to note that while the interpretability mechanisms proposed in these models are technically sophisticated, e.g. counterfactual explanations [31], they often don’t translate well into clinical feedback [16, 15]. Wang et al. [20] proposed a framework that identifies problematic events throughout a surgical trial. They employed a supervised recurrent network to aggregate intermediate Global Rating Scale (GRS) scores into a final score. However, the reliance on more granular and additional labels along with the inherent lack of descriptive qualities in the GRS score are notable drawbacks of their method. Anastasiou et al. [17] adopted a contrastive approach for regressing the GRS score, facilitating a straightforward interpretation of performance relative to a reference. However, deviations from this reference lack translation into actionable feedback or clinical interpretability. A deviation of 10 points from a reference doesn’t allow the trainee to know exactly where he underperformed when the score he receives is an aggregate of 6 distinct surgical skills. Zia et al. [8] investigated how specific segments of the input data, through hand-crafted features, influence OSATS score predictions. While this approach yields more meaningful insights into the variations of performance, they do not offer any actionable information to the clinician. The change in predictions cannot be mapped in a one-to-one fashion to the qualitative description the OSATS give. Similarly, Fawaz et al. [14] extracted performance variations by regressing OSATS scores, whereby GRAD-CAM [27] is used to pinpoint positive and negative contributions to the final output. Nevertheless, both methods fall short of providing clinically comprehensive feedback.
Although interpretable and explainable, these methods often lack the ability to provide actionable insights into trainees’ performance. While they can identify events, they fall short in answering specific questions regarding the nature of their impact. For instance, they may not differentiate whether adverse outcomes stem from inappropriate instrument usage resulting in tissue damage (reflected by a low score in the OSATS category of respect for tissue) or from inefficiencies in movement (indicated by a low score in the OSATS category of time and motion). Having access to such scores throughout a trainee’s trial would allow us to bridge the gap between assessment and training.
In this paper, we propose a novel approach to surgical skill assessment that not only moves beyond the limitations of solely predicting a GRS score but also produces intermediate OSATS scores in a weakly-supervised manner removing the need for additional labels. We introduce a recurrent cross-attention model that leverages kinematic signals to predict six segment-based OSATS scores throughout a surgical trial, providing a more comprehensive understanding of a surgeon’s performance. By correlating these scores with the GRS, our model outperforms existing state-of-the-art (SOTA) models using kinematic signals and offers more granular clinically meaningful insights into surgical proficiency. Our approach facilitates the translation of quantitative predictions into actionable qualitative feedback, which is essential for enhancing automated surgical skill assessment pipelines. We summarise our contributions as follows:
  1. 1.
    A recurrent cross-attention architecture capable of predicting surgical trial level GRS and OSATS scores, while also producing granular segment-level intermediate OSATS scores in a weakly-supervised manner, outperforming existing models using kinematics data,
  2. 2.
    The building blocks for task agnostic models to relate segment-level OSATS predictions to qualitative feedback.
II Methods
We propose a recurrent model called ReCAP, Recursive Cross-Attention for Pseudo-label generation, where segments of kinematic data are processed into intermediate OSATS scores (Fig. 1). Those scores are then averaged into trial-level OSATS predictions. Our multi-task model is trained in an end-to-end fashion on all six OSATS. We assess the model’s performance on the GRS label by aggregating the individual OSATS predicted scores.
II-A Problem Formulation
An input signal $X_{i}\in\mathbb{R}^{D\times T_{i}}$, of feature size D and length $T_{i}$, is divided into equal segments $x^{s}_{i}$ of size $L$ ($x^{s}_{i}\in\mathbb{R}^{D\times L}$): $\{x^{1}_{i},x^{2}_{i},\dots,x^{s}_{i}\}\in X_{i}$ where $S_{i}$ is the total number of segments, i.e. $S_{i}=\frac{T_{i}}{L}$ for given signal index i. For simplicity, in the rest of this paper, we omit $i$ and use s as a subscript to refer to different segments within a trial $i$, i.e. $x^{s}_{i}\to x_{s}$. We fit a function $F$ to map $X$ to the label space $\mathcal{Y}:(F:X\to\vec{y})$:
$\vec{y}=F(x_{1},x_{2},\ldots,x_{s},\ldots x_{S})$ (1)
where $\vec{y}\in\mathcal{Y}$ is a vector composed of all OSATS The GRS is the aggregate of OSATS: $Y=\sum y_{n}$ where $y_{n}$ is the $n^{th}$ OSATS.
Considering clinical practice where the given score is representative of their average performance through the trial i.e.: $y_{n}=\frac{1}{S}\sum y^{s}_{n}$, we rewrite Eq. 1 into Eq. 2, where $f_{n}$ maps a segment to the $n$th OSATS intermediate label ($x_{s}\to y_{s}^{n}$). Note that there is no ground truth for $y_{s}^{n}$ and we learn $\hat{y}_{s}^{n}$ in a weakly-supervised manner.
$y_{n}=\frac{1}{S}\sum_{s=1}^{S}f_{n}(x_{s})$ (2)
II-B Model Overview
Our model processes segments of a kinematic signal recurrently by taking two inputs: the previous hidden state of the recurrent network, $z_{s-1}\in\mathbb{R}^{D\times L}$, and the current segment-level kinematic signal, $x_{s}$ (Fig 1a). The two inputs are fused into the current hidden state, $z_{s}=h(x_{s},z_{s-1})$, through the model backbone $h$ (Fig. 1b). We initialise $z_{0}$ as a zero-filled tensor. Each hidden state is then passed to six classification heads, $c_{n}$, giving $f_{n}(x_{s})=c_{n}(h(x_{s},z_{s-1}))$. The output of our model is a final OSATS, the average of all segment-level OSATS:
$\hat{y_{n}}=\frac{1}{S}\sum\limits_{s=0}^{S}c_{n}[h(x_{s},z_{s-1})]$ (3)
The backbone $h$ is composed of one fusion module (Fig. 1c) similar to Yang et al [10], where previous temporal information is fused with the current input through a series of multi-head self- and cross-attention blocks.
Figure 1: ReCAP Architecture Overview: (a) A kinematic signal is split into segments $x_{s}$ of size $L$ and used as inputs to our backbone $h$ recurrently. The previous state $z_{s-1}$ is also passed as an input and fused to $x_{s}$ to produce $z_{s}$ (c) This fusion is performed by three fusion modules consisting of self- and cross-attention blocks. (b) The current hidden state $z_{s}$ is given as an input to five classification heads $c_{n}$ to predict the respective OSATS scores . $y_{1s}$ to $y_{5s}$ corresponds to the respective OSATS category prediction of Time and Motion, Flow of Operation, Suture/Needle Handling, Respect of Tissue, and Overall Performance at segment position $s$
The classification heads $c_{n}$ are five MLPs classifying the hidden state $z_{s}$ into segment-level OSATS predictions $\hat{y}_{ns}=c_{n}(z_{s})$. Each MLP layer consists of batch normalisation, ReLU activation function, and fully connected layers.
Loss: ReCAP is trained end-to-end using cross-entropy losses. Each cross-entropy loss is applied to the average of the classification head segment predictions $\hat{y_{n}}=\frac{1}{S}\sum\hat{y}_{ns}$ for a given OSATS category label $y_{n}$. An $L2$ penalty term is added to regularize the network and help with generalisation. Our final loss is expressed as:
$\mathcal{L}=\sum_{n=0}^{N}CE(\hat{y_{n}},y_{n})+\lambda*L2$ (4)
TABLE I: Performance comparison of GRS score on JIGSAW for the LOSO cross-validation scheme of models trained on independent and across tasks. K: Kinematic, V Video. *: Results from training on the three domains.
Task & Scheme
Input Method
KT NP SU AT
Spearman’s Correlation Coef (SCC)
[…]
K SMT-DCT-DFT[8] 0.70 0.38 0.64 0.59
K DCT-DFT-ApEn[8] 0.63 0.46 0.75 0.63
[…]
Mean Average Error (MAE)
[…]
II-C Experimental Design
Dataset: We evaluated our model on the JIGSAWS dataset [30]. This dataset consists of video and kinematics data generated by eight clinicians evaluated on three distinct tasks, namely needle passing (NP), suturing (SU), and knot-tying (KT). Altogether, there are 39, 28, and 36 labelled data samples for SU, NP, and KT, respectively. The labels are comprised of six OSATS(1-5) and one GRS(6-35) score the aggregate of all OSATS. It is worth noting that the score 5 only appears in the suturing task. The OSATS include 1) respect for tissue, 2) suture/needle handling, 3) time and motion, 4) flow of operation, 5) overall performance, and 6) quality of the final product. In this study, we only use kinematic data.
Cross-validation: Following the JIGSAW cross-validation framework, we evaluate our method using Leave-One-Supertrial-Out (LOSO) whereby the i-th trial performed by the surgeons are left out as the validation set. Another cross-validation scheme Leave-One-User-Out (LOUO) is not considered in this work. The lack of literature reporting OSATS performance on this scheme and very imbalanced training/validation folds were reasons enough. Recent literature [32, 17] on the Jigsaw dataset seems to indicate that no testing fold is used to assess performance.
Performance Metric and Reporting: Similar to relevant work [8, 6, 7, 28], we evaluate our method using Spearman’s Correlation Coefficient (SCC) $\rho$ to compare the predicted ranked GRS score with the ground truth (Eq. 5). We report SCC averaged across folds for the LOSO cross-validation scheme under the same training parameters and at the same training epoch. The intermediate OSATS are averaged into a signal-level OSATS after processing the whole kinematic sample. The final six OSATS are then summed to give a video-level GRS score. Note that the predicted GRS score, $\hat{Y}=\sum_{n=0}^{6}\hat{y_{n}}$, is only used to assess model performance, but is not directly learned from our model:
$\rho=1-\frac{6\sum d_{i}^{2}}{n(n^{2}-1)}=1-\frac{6\sum{(Y-\sum_{n=0}^{6}\hat{y_{n}})}^{2}}{n(n^{2}-1)}$ (5)
The mean average error (MAE) is also reported at the best-performing epoch for this specific metric. To our knowledge, Benmansour et al [32] are the only recent work reporting OSATS-specific performance under the LOSO validation scheme. OSATS performance is reported at the same epoch used to report GRS performance under the same training parameters. OSATS SCC was averaged across 10 epochs to mitigate the intrinsic variability coming from the short range of labels. OSATS are learned and regressed by our model.
Data Augmentation: Two augmentation techniques were added to the kinematic signals to improve generalisation: 1) Gaussian noise based on the standard deviation of the signal, and 2) flipping, i.e. reversing the signal. Augmentations were done at a rate of $50\%$. Label smoothing was also performed at 30%.
Implementation: Our model is trained with the Adam optimizer for 5000 epochs with a learning rate of $10^{-6}$. Kinematic data was pre-processed as suggested by De Iturrate et al. [9] by normalising across time and feature dimensions. The kinematics from the slave and master device were used ($D=76$). All experiments were seeded for reproducibility purposes. The sequence length of 75 was chosen. It corresponds to 2.5s in time(force data acquired at 30hz) and is consistent with the minimal time required for our clinician to rate a gesture. A lambda of 0.01 for L2 regularization and a batch size of 25 was used. In line with existing literature, design decisions and hyperparameter adjustments were experimentally conducted using the averaged cross-validation test fold [17]. Hyperparameters were kept the same across all tasks. ReCAP was implemented in Pytroch and trained on an Nvidia A100 GPU.
Validation of Model Behaviour: To validate our model’s ability to generate interim OSATS scores, we asked a consultant surgeon in endoscopic interventions to agree or disagree with the model’s intermediate predictions. Every 75 frames were assigned a generated pseudo-label. The label is shown on the screen during viewing. Similar to Wang et al.’s framework [20], the surgeon was aware of the ground truth assigned to the video. The predicted OSATS overall performance score was divided into three categories: poor (1-2), average (3), and good (4-5). We randomly introduce noise to the model’s predictions to mitigate potential bias without informing the surgeon. We then present these predictions and capture agreement or disagreement at the segment level when playing the video sequentially.
TABLE II: Perofrmance for OSATS scores, where the $\rho Osats$ is the average across the 6 scores under LOSO scheme. *: results from training across the 3 tasks.
KT NP SU AT
Apen[8] 0.66 0.45 0.59 0.57
[…]
III Results
TABLE III: Performance for OSATS scores, where the $\rho Osats$ are reported for RT: Respect for tissue. TM: Time and Motion. OP: Overall Performance under the LOSO in the KT task.
CNN + Bilstm     [32] ReCAP
[…]
RT 0.83 0.49 0.46 0.92/0.78 0.75/0.43 0.78/0.52 0.56
TM 0.87 0.85 0.68 0.95/0.8 0.91/0.72 0.84/0.60 0.62
OP 0.89 0.58 0.71 0.9/0.79 0.42/0.23 0.69/0.5 0.65
SNH 0.82 0.79 0.75 0.84/0.61 0.91/0.69 0.88/0.78 0.65
FO 0.76 0.58 0.62 0.78/0.63 0.66/0.45 0.89/0.66 0.64
QFP 0.75 0.31 0.67 0.85/0.59 0.56/0.22 0.91/0.64 0.62
Mean 0.82 0.60 0.65 0.87/0.70 0.70/0.46 0.83/0.62 0.62
We report the performance of our model against previous work that uses kinematic data or video data and report GRS performance (Table II-B). Although we don’t regress the GRS’s most recent work only report on it. To allow for easier comparison we use the GRS as a performance proxy. The model outperforms all methods using kinematic data and achieves competitive performance against models using video (Table II-B). When looking at the performance of our model in predicting OSATS scores under the LOSO validation scheme, we underperform only in NP (Table II). The CNN+Bilstm [32] only reports the best-performing fold. We also report the average across the 5 folds. We only underperform for the OSATS of Overall Performance in needle passing.
We performed two ablation studies to understand the contribution of each element of the model on the performance. The introduced guassian noise and flipping had very little effect on the performance of the model. However the the flipping does allow for the model to be time invariant. We see that the pseudo-label drastically improves performance, especially for the two tasks, NP and SU, with the most class imbalance [29].
TABLE IV: Ablation of ReCAP components for GRS under LOSO scheme.
KT NP SU
ReCAP no augmentation 0.86 0.85 0.83
[…]
To validate the weakly supervised outputs, 9 videos were reviewed by a consultant surgeon, where each 75 frames (the segment length) had an assigned OSATS pseudo-label. The selected videos regrouped three levels of expertise (novice, intermediate, expert) across the three tasks. Two of those videos were shown with randomly generated predictions. We found that the clinician agreed 69% of the time when shown random noise while agreeing 77% of the time when shown our model’s predictions. A one-tailed binomial test between the two distributions indicates a statistically significant difference between the agreements (p=0.006).
Figure 2: Variations of OSATS scores for a Knot Tying task by a self-proclaimed expert. Qualitative descriptions are taken from [19]. In this example, the senior clinician disagreed with the model’s intermediate scores in 3 instances.
IV Discussion
IV-A Performance
To the best of our knowledge, we are the only work to report performance metrics on the JigSaw Dataset in a task-agnostic manner. Recent studies have focused on video-centric models, which don’t generalise well outside of their training domain, whereas pure kinematic, a task-agnostic modality, has seen decreased interest. This research showcases that there is still substantial room for improvement in the kinematic domain on the JigSaw dataset. This work showcased competitive performance on the GRS. It doesn’t beat SOA in the video domain, certainly due to overfitting. The ratio between input features and model parameters will always lead to overfitting in deep-learning applications (238644 / 1440). On the OSATS, our model performs better than most, and we thoroughly report on our performance. We do see that for Overall Performance in Needle Passing and Quality of the Final Product across the three tasks, the model doesn’t perform as well. We argue that this is because while kinematic data is an extremely powerful modality it cannot capture all the nuances in performance. It was shown that the quality-of-final-product by Kasa et al. [21] is an image/video-centric score. Furthermore, it is also worth noting that similar kinematic profiles might end up with very different performance scores. In needle passing if a clinician misgauges the depth, he will fail the task, and receive a low Overall Performance score which wouldn’t necessarily be translated in their kinematic profile. As highlighted by Lefor et al. [29], the JIGSAW dataset is very imbalanced and inverse correlation among subjects can be observed. Furthermore using Speaman’s correlation coefficient when there are sometimes only 3 test samples could be misleading. Consequently, developing tools based on this dataset may not seamlessly translate to real-world practice. We argue that our competitive results, despite a simple architecture, are because our problem formulation allows the model to attend segments of the input independently, while retaining some temporal context, allowing for much more flexible behaviour and sparse solutions.
IV-B Proposed Method
Through the ablations, we showcase that using the pseudo-label in the loss boosts the model’s performance. We hypothesise that including intermediate predictions prevents overfitting and acts as a regularization component. Furthermore this also fits within how raters would look at performance, where each previous segment affects their decision. However the way the loss is built wouldn’t allow for the model to understand when a catastrophic event would occur, which would always result in a systematic low performance score. Allowing the model to assign a weight to the segments should adress this issue.
IV-C Pseudo-Labels
The contribution of this loss is two-fold, improving performance and generating pseudo-labels. We extended the generation of pseudo-labels to insights into the user’s performance. We can visualize our pipeline results in Fig. 2. As the introduction emphasised, mapping performance to actionable feedback is necessary for clinicians’ understanding and the first steps toward clinical translation of such models. It is also worth noting the recurrent nature of the model allows for its online application. However, the validation of those pseudo-labels has serious limitations. Distinguishing between good/ average and average/poor levels at such a fine-grained scale presents significant complexity for any rater. This is obvious when looking at the agreement rate of the surgeon with random noise. The expected agreement rate should’ve been around 33%. Additionally, variability among raters is commonly encountered in various fields and tasks, making the extraction of ground truth complex. In an ideal scenario, having more data and involving multiple raters could help mitigate this variability. In practice, this is not feasible, and methods to extract labels in a weakly-supervised might be a more promising approach. The significant difference between random noise and the model’s prediction showcases a promising first step towards that goal, for both performance gains and interpretability.
V Conclusion and Future Work
In this work, we adopt a different take on the problem formulation for skill assessment which can be expanded to more complex recurrent architectures and other fields of skill assessment. Through our competitive results on the JIGSAW dataset, we demonstrate the feasibility of this approach, while providing more granular performance insights for skill assessment. In future work, our focus will be on enhancing the robustness of validation, incorporating other time series data that an operating room could generate [33], and extending our approach to datasets featuring longer and more complex tasks. Surgical tasks/interventions lasting multiple hours are very demanding to annotate. Semi-supervised methods that extract intermediate labels, at the gesture, step, and phase levels could allow for more automated and more granular assessment of surgeons’ performance.